# Compact manifolds of dimension $n\geq 12$ with positive isotropic curvature

### v21

@@ 5 Appendix: gluing and strengthening of $\varepsilon$-necks
Proof.  Let $N_{0}$ be the $\varepsilon$-neck in $M$ centered at $\gamma(0)$, and $S$ be the central cross section of the neck $N_{0}$.
@@
Proof of Claim. We argue by contradiction. Suppose the claim is not true. Choose $s_{0}>0$ such that $\gamma(s_{0})\in N_{0}$, and a point $x_{1}\in N_{0}\setminus S$ such that $x_{1}$ and $\gamma(s_{0})$ lie in different components of $N_{0}\setminus S$. Choose a path $\alpha:[0,1]\rightarrow M\setminus S$ with $\alpha(0)=\gamma(s_{0})$ and $\alpha(1)=x_{1}$. Note that there exists $s_{1}\in(0,1)$ such that for any $s\in[s_{1},1)$, the intersection of the (image of the) path $\alpha$ with the central cross section of the $\varepsilon$-neck centered at $\gamma(s)$ is empty, since $R(\gamma(s))\rightarrow\infty$ as $s\rightarrow 1$ but the image of $\alpha$ is compact. Denote the central cross section of the $\varepsilon$-neck centered at $\gamma(s_{1})$ by $S_{1}$. We may assume that $S\cap S_{1}=\emptyset$ by choosing $s_{1}$ with $1-s_{1}$ small. By Proposition 5.4 the hypersurface $S$ and $S_{1}$ cobound a compact domain in $M$, denoted by $\Omega$, which is diffeomorphic to $\mathbb{S}^{n-1}\times[0,1]$ and contains $\gamma(s_{0})$ but does not contain $x_{1}$. As the intersection of the (image of the) path $\alpha$ with $\partial\Omega$ is empty, and $\alpha(0)=\gamma(s_{0})\in\Omega$, it follows that the image of $\alpha$ is contained in $\Omega$. But on the other hand $\alpha(1)=x_{1}\notin\Omega$. A contradiction. $\Box$
Now we remove from $M$ the component of $M\setminus S$ which does not contain $\gamma((0,1))$, and get a manifold, denoted by $Y$, which contains $\gamma([0,1))$ and has $S$ as its boundary. Glue smoothly a $n$-disk $D_{0}$ to $Y$ along their boundaries, and denote the manifold thus obtained by $Z$. Note that $Z$ is a connected, noncompact, smooth manifold without boundary. Using Proposition 5.4 and arguing as in the proof of Lemma 3.2.8 in [6] we see that $Z=\cup_{i=0}^{\infty}D_{i}$, where each $D_{i}$ is diffeomorphic to $D^{n}$, and $D_{i}\subset\text{Int}\hskip 1.42262ptD_{i+1}$. Now using the Cerf-Palais disk theorem (see for example Theorem 3.1 in Chapter 8 of [49] or Theorem 2.1 on p. 197 of [66]) one can show that $Z$ is diffeomorphic to $\mathbb{R}^{n}$ (see p. 126 of [66]). By the Cerf-Palais disk theorem again $Z\setminus D_{0}$ is diffeomorphic to $\mathbb{S}^{n-1}\times(0,1)$, and the desired result follows. $\Box$
@@
Remark. Consider the identity map $\text{id}:(\mathbb{S}^{n-1}\times(a,b),g_{\text{cly}})\rightarrow(\mathbb{S}^{n-1}\times(a,b),cg_{\text{cly}})$, where $c$ is an arbitrary positive constant. If we want the map id to be a Hamilton’s canonical parametrization, then the exponent appearing in the integrand on the RHS of the equality (5.1) should be $n$ instead of $n-1$. Compare Remark 3.10 (iii) in [54].
The following lemma is adapted from Theorem C2.2 and Corollary C2.3 in [48], and is a preparation for Proposition 5.7.
Lemma 5.6.
@@
Proof.   Let $0<\varepsilon<\frac{1}{10^{3}n}$ be given. Suppose that $N$ is an $\eta$-neck with scaling factor $Q$ in $(\mathcal{O},g)$ given by a diffeomorphism $\psi:\mathbb{S}^{n-1}/\Gamma\times(-\eta^{-1},\eta^{-1})\rightarrow N$ for some $\eta>0$. Consider the Riemannian universal cover $\pi:\tilde{N}\rightarrow N$ of the neck $N$, which has an induced $\eta$-neck structure given by a lift $\tilde{\psi}:\mathbb{S}^{n-1}\times(-\eta^{-1},\eta^{-1})\rightarrow\widetilde{N}$ of $\psi$. Let $\widetilde{\Gamma}$ be the deck transformation group of the Riemannian covering $\pi:\tilde{N}\rightarrow N$. We have $\widetilde{\Gamma}\cong\Gamma$. Note that for any $z\in(-\eta^{-1},\eta^{-1})$, $\tilde{\psi}(\mathbb{S}^{n-1}\times\{z\})$ is $\widetilde{\Gamma}$-invariant. By (the proof of) Theorem C2.2 and Corollary C2.3 in [48], there exists a positive number $\tilde{\varepsilon}_{1}=\tilde{\varepsilon}_{1}(\varepsilon)$ with the following property. If $0<\eta\leq\tilde{\varepsilon}_{1}$, there exists a Hamilton’s canonical parametrization $\tilde{\phi}:\mathbb{S}^{n-1}\times(a,b)\rightarrow\widetilde{\mathcal{T}}\subset\widetilde{N}$ which gives an $\varepsilon$-tube $\widetilde{\mathcal{T}}$ in $\widetilde{N}$ with $\tilde{\psi}(\mathbb{S}^{n-1}\times(-\eta^{-1}+1,\eta^{-1}-1))\subset\widetilde{\mathcal{T}}$, in the sense that for any $(c,d)\subset(a,b)$ with $d-c=2\varepsilon^{-1}$, $\tilde{\phi}|_{\mathbb{S}^{n-1}\times(c,d)}$ gives an $\varepsilon$-neck structure (up to a translation of the parameter) for its image, such that for each $s\in(a,b)$, $\tilde{\phi}(\mathbb{S}^{n-1}\times\{s\})$ is a constant mean curvature hypersurface which is close (with error controlled in terms of $\varepsilon$) to horizontal w.r.t. the $\eta$-neck structure of $\widetilde{N}$. The closeness of the constant mean curvature hypersurface $\tilde{\phi}(\mathbb{S}^{n-1}\times\{s\})$ to the corresponding horizontal hypersurface (w.r.t. the $\eta$-neck structure on $\widetilde{N}$) can be described more precisely as follows. Given any $s\in(a,b)$, and $\tilde{x}_{0}\in\tilde{\phi}(\mathbb{S}^{n-1}\times\{s\})$, let $\widetilde{\Sigma}_{0}$ be the cross section of the neck $\widetilde{N}$ passing through $\tilde{x}_{0}$, then we have $\tilde{\phi}(\mathbb{S}^{n-1}\times\{s\})=\{\exp_{\tilde{x}}(Q^{-1/2}\tilde{f}(\tilde{x})\tilde{\nu}(\tilde{x}))\hskip 2.84526pt|\hskip 2.84526pt\tilde{x}\in\widetilde{\Sigma}_{0}\}$, where $\exp$ is the exponential map of the Riemannian manifold $\widetilde{N}$, $\tilde{\nu}$ is a unit normal vector field on $\widetilde{\Sigma}_{0}$, and $\tilde{f}$ is certain smooth function on $\widetilde{\Sigma}_{0}$ with $\tilde{f}(\tilde{x}_{0})=0$ and $||\tilde{f}||_{C^{4,\frac{1}{2}}}<\frac{1}{2}\varepsilon$. Here the $C^{4,\frac{1}{2}}$-norm is taken w.r.t. the round metric on $\widetilde{\Sigma}_{0}$ with scalar curvature 1. See also the proof of Proposition D.1 in version 1 of [22] on arXiv. We take $\tilde{\phi}$ to be maximal in the sense that $\widetilde{\mathcal{T}}$ cannot be properly included in the image of any Hamilton’s canonical parametrization which gives an $\varepsilon$-tube in $\widetilde{N}$.
@@
Let $s\in(a,b)$, $\tilde{x}_{0}\in\tilde{\phi}(\mathbb{S}^{n-1}\times\{s\})$, and $\widetilde{\Sigma}_{0}$ be the cross section of the neck $\widetilde{N}$ passing through $\tilde{x}_{0}$ as above. Let $x_{0}=\pi(\tilde{x}_{0})$, and $\Sigma_{0}=\pi(\widetilde{\Sigma}_{0})$. As both the ambient Riemannian manifold $\widetilde{N}$ and the hypersurfaces $\widetilde{\Sigma}_{0}$ and $\tilde{\phi}(\mathbb{S}^{n-1}\times\{s\})$ are $\widetilde{\Gamma}$-invariant, the function $\tilde{f}$ above must also be $\widetilde{\Gamma}$-invariant, so it descends to a smooth function $f$ on $\Sigma_{0}$ with $f(x_{0})=0$ and $||f||_{C^{4,\frac{1}{2}}}<\frac{1}{2}\varepsilon$ such that $\phi(\mathbb{S}^{n-1}/\widehat{\Gamma}\times\{s\})=\{\exp_{x}(Q^{-1/2}f(x)\nu(x))\hskip 2.84526pt|\hskip 2.84526ptx\in\Sigma_{0}\}$. Here the $C^{4,\frac{1}{2}}$-norm is taken w.r.t. the round metric on $\Sigma_{0}$ with scalar curvature 1, $\exp$ is the exponential map of the Riemannian manifold $N$, and $\nu$ is a unit normal vector field on $\Sigma_{0}$. Now we see that $\phi(\mathbb{S}^{n-1}/\widehat{\Gamma}\times\{s\})$ is diffeomorphic to $\Sigma_{0}$. Note that $\Sigma_{0}$ is diffeomorphic to $\mathbb{S}^{n-1}/\Gamma$. It follows that $\widehat{\Gamma}$ is conjugate to $\Gamma$ in $O(n)$, and we can view the Hamilton’s canonical parametrization $\phi:\mathbb{S}^{n-1}/\widehat{\Gamma}\times(a,b)\rightarrow\mathcal{T}$ as a map from $\mathbb{S}^{n-1}/\Gamma\times(a,b)$ to $\mathcal{T}$ which is still a Hamilton’s canonical parametrization. $\Box$
@@
The following result is based on Lemma 5.6 above and Theorem C2.4 in [48]. For relevant discussion in dimension $n=4$ see Appendix B in [50] and Proposition 2.3 in [51].
Proposition 5.7.
@@
(cf. Theorem C2.4 in [48]) Let $n\geq 4$. There exists a positive constant $\varepsilon_{3}=\varepsilon_{3}(n)\leq\varepsilon_{2}$, where $\varepsilon_{2}$ is as in Lemma 5.6, with the following property. Let $0<\varepsilon\leq\varepsilon_{3}$, and $\tilde{\varepsilon}_{1}=\tilde{\varepsilon}_{1}(\varepsilon)$, where $\tilde{\varepsilon}_{1}(\cdot)$ is as in Lemma 5.6. Let $(\mathcal{O},g)$ be a Riemannian orbifold of dimension $n$ with at most isolated singularities, $0<\eta\leq\tilde{\varepsilon}_{1}$, and $N_{i}$ be an $\eta$-neck centered at a point $x_{i}$ in $\mathcal{O}$ given by a diffeomorphism $\psi_{i}:\mathbb{S}^{n-1}/\Gamma_{i}\times(-\eta^{-1},\eta^{-1})\rightarrow N_{i}$, $i=1,2$, such that $\psi_{1}(\mathbb{S}^{n-1}/\Gamma_{1}\times(-\eta^{-1}+2,\eta^{-1}-2))\cap\psi_{2}(\mathbb{S}^{n-1}/\Gamma_{2}\times(-\eta^{-1}+1,\eta^{-1}-1))\neq\emptyset$. Then $\Gamma_{1}$ and $\Gamma_{2}$ are conjugate in $O(n)$. Moreover, if $\psi_{1}(\mathbb{S}^{n-1}/\Gamma_{1}\times(-\eta^{-1},-\eta^{-1}+0.01))\cap N_{2}=\emptyset$ and $\psi_{2}(\mathbb{S}^{n-1}/\Gamma_{2}\times(\eta^{-1}-0.01,\eta^{-1}))\cap N_{1}=\emptyset$, there exists a Hamilton’s canonical parametrization $\phi:\mathbb{S}^{n-1}/\Gamma_{1}\times(a,b)\rightarrow\mathcal{T}\subset N_{1}\cup N_{2}$ which gives an $\varepsilon$-tube $\mathcal{T}$ in $N_{1}\cup N_{2}$ with $\psi_{1}(\mathbb{S}^{n-1}/\Gamma_{1}\times(-\eta^{-1}+1,\eta^{-1}))\cup\psi_{2}(\mathbb{S}^{n-1}/\Gamma_{2}\times(-\eta^{-1},\eta^{-1}-1))\subset\mathcal{T}$.
@@
Proof.   Given $0<\varepsilon\leq\varepsilon_{2}$ and $\eta$-neck $N_{i}$ ($i=1,2$) in $(\mathcal{O},g)$ with $\eta\leq\tilde{\varepsilon}_{1}(\varepsilon)$, by Lemma 5.6, there exists a Hamilton’s canonical parametrization $\phi_{i}:\mathbb{S}^{n-1}/\Gamma_{i}\times(a_{i},b_{i})\rightarrow\mathcal{T}_{i}\subset N_{i}$ which gives an $\varepsilon$-tube $\mathcal{T}_{i}$ in $N_{i}$ with $\psi_{i}(\mathbb{S}^{n-1}/\Gamma_{i}\times(-\eta^{-1}+1,\eta^{-1}-1))\subset\mathcal{T}_{i}$. We can take $\phi_{i}$ to be maximal in the sense that $\mathcal{T}_{i}$ cannot be properly included in the image of any Hamilton’s canonical parametrization which gives an $\varepsilon$-tube in $N_{i}$, $i=1,2$. Let $\Phi_{i}=\phi_{i}\circ\pi_{i}:\mathbb{S}^{n-1}\times(a_{i},b_{i})\rightarrow\mathcal{T}_{i}\subset N_{i}$ be Hamilton’s canonical uniformization of $\mathcal{T}_{i}$, where $\pi_{i}:\mathbb{S}^{n-1}\times(a_{i},b_{i})\rightarrow\mathbb{S}^{n-1}/\Gamma_{i}\times(a_{i},b_{i})$ is the natural projection. Choose a point $x_{0}\in\psi_{1}(\mathbb{S}^{n-1}/\Gamma_{1}\times(-\eta^{-1}+2,\eta^{-1}-2))\cap\psi_{2}(\mathbb{S}^{n-1}/\Gamma_{2}\times(-\eta^{-1}+1,\eta^{-1}-1))$. Then there exist $s_{1}\in(a_{1},b_{1})$ away from $a_{1}$ and $b_{1}$, and $s_{2}\in(a_{2},b_{2})$, such that $x_{0}\in\Phi_{1}(\mathbb{S}^{n-1}\times\{s_{1}\})\cap\Phi_{2}(\mathbb{S}^{n-1}\times\{s_{2}\})$. If $\varepsilon_{3}=\varepsilon_{3}(n)\leq\varepsilon_{2}$ is sufficiently small and $0<\varepsilon\leq\varepsilon_{3}$, by the proof of Theorem C2.4 in [48], which uses Lemma C2.1 in [48], there exists an isometry of the round cylinder $\mathbb{S}^{n-1}\times\mathbb{R}$, say $G$, such that $G(\mathbb{S}^{n-1}\times\{s_{2}\})=\mathbb{S}^{n-1}\times\{s_{1}\}$ and $\Phi_{1}G|_{\mathbb{S}^{n-1}\times\{s_{2}\}}=\Phi_{2}|_{\mathbb{S}^{n-1}\times\{s_{2}\}}$. So $\phi_{1}(\mathbb{S}^{n-1}/\Gamma_{1}\times\{s_{1}\})=\phi_{2}(\mathbb{S}^{n-1}/\Gamma_{2}\times\{s_{2}\})$. It follows that $\Gamma_{1}$ and $\Gamma_{2}$ are conjugate in $O(n)$.
@@
As in Section 72 of [55], there exist numbers $B$ and $B^{\prime}$ with $\frac{1}{10}<B^{\prime}<B$, and a smooth convex function $\varphi$ defined on $(-\infty,B]$ with $\varphi|_{(-\infty,0]}=0$ and $\varphi(z)=ce^{-1/z}$ for $z\in(0,\frac{1}{10}]$, where $c<\frac{c_{n}}{2}$ is a positive constant, such that $\varphi$ coincides with $\varphi_{k}$ on $[B^{\prime},B]$ for some $k>0$, and the metric $e^{-2\varphi}g_{\text{cly}}|_{\mathbb{S}^{n-1}\times(-\infty,B]}$ has nonnegative curvature operator. Here, we use the fact that if $c<\frac{c_{n}}{2}$,

### v22

@@ Appendix A Appendix: gluing and strengthening of $\varepsilon$-necks
Proof.  Let $N_{0}$ be the $\varepsilon$-neck in $M$ centered at $\gamma(0)$, and $S$ be the central cross-section of the neck $N_{0}$.
@@
Proof of Claim. We argue by contradiction. Suppose the claim is not true. Choose $s_{0}>0$ such that $\gamma(s_{0})\in N_{0}$, and a point $x_{1}\in N_{0}\setminus S$ such that $x_{1}$ and $\gamma(s_{0})$ lie in different components of $N_{0}\setminus S$. Choose a path $\alpha:[0,1]\rightarrow M\setminus S$ with $\alpha(0)=\gamma(s_{0})$ and $\alpha(1)=x_{1}$. Note that there exists $s_{1}\in(0,1)$ such that for any $s\in[s_{1},1)$, the intersection of the (image of the) path $\alpha$ with the central cross section of the $\varepsilon$-neck centered at $\gamma(s)$ is empty, since $R(\gamma(s))\rightarrow\infty$ as $s\rightarrow 1$ but the image of $\alpha$ is compact. Denote the central cross section of the $\varepsilon$-neck centered at $\gamma(s_{1})$ by $S_{1}$. We may assume that $S\cap S_{1}=\emptyset$ by choosing $s_{1}$ with $1-s_{1}$ small. By Proposition A.4 and Lemma 3.1 the hypersurface $S$ and $S_{1}$ cobound a compact domain in $M$, denoted by $\Omega$, which is diffeomorphic to $\mathbb{S}^{n-1}\times[0,1]$ and contains $\gamma(s_{0})$ but does not contain $x_{1}$. As the intersection of the (image of the) path $\alpha$ with $\partial\Omega$ is empty, and $\alpha(0)=\gamma(s_{0})\in\Omega$, it follows that the image of $\alpha$ is contained in $\Omega$. But on the other hand $\alpha(1)=x_{1}\notin\Omega$. A contradiction. $\Box$
Now we remove from $M$ the component of $M\setminus S$ which does not contain $\gamma((0,1))$, and get a manifold, denoted by $Y$, which contains $\gamma([0,1))$ and has $S$ as its boundary. Glue smoothly a $n$-disk $D_{0}$ to $Y$ along their boundaries, and denote the manifold thus obtained by $Z$. Note that $Z$ is a connected, noncompact, smooth manifold without boundary. Using Proposition A.4 and arguing as in the proof of Lemma 3.2.8 in [6] we see that $Z=\cup_{i=0}^{\infty}D_{i}$, where each $D_{i}$ is diffeomorphic to $D^{n}$, and $D_{i}\subset\text{Int}\hskip 1.42262ptD_{i+1}$. Now using the Cerf-Palais disk theorem (see for example Theorem 3.1 in Chapter 8 of [52] or Theorem 2.1 on p. 197 of [69]) one can show that $Z$ is diffeomorphic to $\mathbb{R}^{n}$ (see p. 126 of [69]). By the Cerf-Palais disk theorem again $Z\setminus D_{0}$ is diffeomorphic to $\mathbb{S}^{n-1}\times(0,1)$, and the desired result follows. $\Box$
@@
Remark. Consider the identity map $\text{id}:(\mathbb{S}^{n-1}\times(a,b),g_{\text{cly}})\rightarrow(\mathbb{S}^{n-1}\times(a,b),cg_{\text{cly}})$, where $c$ is an arbitrary positive constant. If we want the map id to be a Hamilton’s canonical parametrization, then the exponent appearing in the integrand on the RHS of the equality (A.1) should be $n$ instead of $n-1$. Compare Remark 3.10 (iii) in [57].
The following lemma is adapted from Theorem C2.2 and Corollary C2.3 in [51], and is a preparation for Proposition A.7.
Lemma A.6.
@@
Proof.   Let $0<\varepsilon<\frac{1}{10^{4}n}$ be given. Suppose that $N$ is an $\eta$-neck with scaling factor $Q$ in $(\mathcal{O},g)$ given by a diffeomorphism $\psi:\mathbb{S}^{n-1}/\Gamma\times(-\eta^{-1},\eta^{-1})\rightarrow N$ for some $\eta>0$. Consider the Riemannian universal cover $\pi:\tilde{N}\rightarrow N$ of the neck $N$, which has an induced $\eta$-neck structure given by a lift $\tilde{\psi}:\mathbb{S}^{n-1}\times(-\eta^{-1},\eta^{-1})\rightarrow\widetilde{N}$ of $\psi$. Let $\widetilde{\Gamma}$ be the deck transformation group of the Riemannian covering $\pi:\tilde{N}\rightarrow N$. We have $\widetilde{\Gamma}\cong\Gamma$. Note that for any $z\in(-\eta^{-1},\eta^{-1})$, $\tilde{\psi}(\mathbb{S}^{n-1}\times\{z\})$ is $\widetilde{\Gamma}$-invariant. By (the proof of) Theorem C2.2 and Corollary C2.3 in [51], there exists a positive number $\tilde{\varepsilon}_{1}=\tilde{\varepsilon}_{1}(\varepsilon)$ with the following property. If $0<\eta\leq\tilde{\varepsilon}_{1}$, there exists a Hamilton’s canonical parametrization $\tilde{\phi}:\mathbb{S}^{n-1}\times(a,b)\rightarrow\widetilde{\mathcal{T}}\subset\widetilde{N}$ which gives an $\varepsilon$-tube $\widetilde{\mathcal{T}}$ in $\widetilde{N}$ with $\tilde{\psi}(\mathbb{S}^{n-1}\times(-\eta^{-1}+1,\eta^{-1}-1))\subset\widetilde{\mathcal{T}}$, in the sense that for any $(c,d)\subset(a,b)$ with $d-c=2\varepsilon^{-1}$, $\tilde{\phi}|_{\mathbb{S}^{n-1}\times(c,d)}$ gives an $\varepsilon$-neck structure (up to a translation of the parameter) for its image, such that for each $s\in(a,b)$, $\tilde{\phi}(\mathbb{S}^{n-1}\times\{s\})$ is a constant mean curvature hypersurface which is close (with error controlled in terms of $\varepsilon$) to horizontal w.r.t. the $\eta$-neck structure of $\widetilde{N}$. The closeness of the constant mean curvature hypersurface $\tilde{\phi}(\mathbb{S}^{n-1}\times\{s\})$ to the corresponding horizontal hypersurface (w.r.t. the $\eta$-neck structure on $\widetilde{N}$) can be described more precisely as follows. Given any $s\in(a,b)$, and $\tilde{x}_{0}\in\tilde{\phi}(\mathbb{S}^{n-1}\times\{s\})$, let $\widetilde{\Sigma}_{0}$ be the cross section of the neck $\widetilde{N}$ passing through $\tilde{x}_{0}$, then we have $\tilde{\phi}(\mathbb{S}^{n-1}\times\{s\})=\{\exp_{\tilde{x}}(Q^{-1/2}\tilde{f}(\tilde{x})\tilde{\nu}(\tilde{x}))\hskip 2.84526pt|\hskip 2.84526pt\tilde{x}\in\widetilde{\Sigma}_{0}\}$, where $\exp$ is the exponential map of the Riemannian manifold $\widetilde{N}$, $\tilde{\nu}$ is a unit normal vector field on $\widetilde{\Sigma}_{0}$, and $\tilde{f}$ is certain smooth function on $\widetilde{\Sigma}_{0}$ with $\tilde{f}(\tilde{x}_{0})=0$ and $||\tilde{f}||_{C^{4,\frac{1}{2}}}<\frac{1}{2}\varepsilon$. Here the $C^{4,\frac{1}{2}}$-norm is taken w.r.t. the round metric on $\widetilde{\Sigma}_{0}$ with scalar curvature 1. See also the proof of Proposition D.1 in version 1 of [23] on arXiv. We take $\tilde{\phi}$ to be maximal in the sense that $\widetilde{\mathcal{T}}$ cannot be properly included in the image of any Hamilton’s canonical parametrization which gives an $\varepsilon$-tube in $\widetilde{N}$.
@@
Let $s\in(a,b)$, $\tilde{x}_{0}\in\tilde{\phi}(\mathbb{S}^{n-1}\times\{s\})$, and $\widetilde{\Sigma}_{0}$ be the cross section of the neck $\widetilde{N}$ passing through $\tilde{x}_{0}$ as above. Let $x_{0}=\pi(\tilde{x}_{0})$, and $\Sigma_{0}=\pi(\widetilde{\Sigma}_{0})$. As both the ambient Riemannian manifold $\widetilde{N}$ and the hypersurfaces $\widetilde{\Sigma}_{0}$ and $\tilde{\phi}(\mathbb{S}^{n-1}\times\{s\})$ are $\widetilde{\Gamma}$-invariant, the function $\tilde{f}$ above must also be $\widetilde{\Gamma}$-invariant, so it descends to a smooth function $f$ on $\Sigma_{0}$ with $f(x_{0})=0$ and $||f||_{C^{4,\frac{1}{2}}}<\frac{1}{2}\varepsilon$ such that $\phi(\mathbb{S}^{n-1}/\widehat{\Gamma}\times\{s\})=\{\exp_{x}(Q^{-1/2}f(x)\nu(x))\hskip 2.84526pt|\hskip 2.84526ptx\in\Sigma_{0}\}$. Here the $C^{4,\frac{1}{2}}$-norm is taken w.r.t. the round metric on $\Sigma_{0}$ with scalar curvature 1, $\exp$ is the exponential map of the Riemannian manifold $(N,g)$, and $\nu$ is a unit normal vector field on $\Sigma_{0}$. Now we see that $\phi(\mathbb{S}^{n-1}/\widehat{\Gamma}\times\{s\})$ is diffeomorphic to $\Sigma_{0}$. Note that $\Sigma_{0}$ is diffeomorphic to $\mathbb{S}^{n-1}/\Gamma$. It follows that $\widehat{\Gamma}$ is conjugate to $\Gamma$ in $O(n)$, and we can view the Hamilton’s canonical parametrization $\phi:\mathbb{S}^{n-1}/\widehat{\Gamma}\times(a,b)\rightarrow\mathcal{T}$ as a map from $\mathbb{S}^{n-1}/\Gamma\times(a,b)$ to $\mathcal{T}$ which is still a Hamilton’s canonical parametrization. $\Box$
@@
The following result is based on Lemma A.6 above and Theorem C2.4 in [51]. For relevant discussion in dimension $n=4$ see Appendix B in [53] and Proposition 2.3 in [54].
Proposition A.7.
@@
(cf. Theorem C2.4 in [51]) Let $n\geq 4$. There exists a positive constant $\varepsilon_{3}=\varepsilon_{3}(n)\leq\varepsilon_{2}$, where $\varepsilon_{2}$ is as in Lemma A.6, with the following property. Let $0<\varepsilon\leq\varepsilon_{3}$, and $\tilde{\varepsilon}_{1}=\tilde{\varepsilon}_{1}(\varepsilon)$, where $\tilde{\varepsilon}_{1}(\cdot)$ is as in Lemma A.6. Let $(\mathcal{O},g)$ be a Riemannian orbifold of dimension $n$ with at most isolated singularities, $0<\eta\leq\tilde{\varepsilon}_{1}$, and $N_{i}$ be an $\eta$-neck centered at a point $x_{i}$ in $\mathcal{O}$ given by a diffeomorphism $\psi_{i}:\mathbb{S}^{n-1}/\Gamma_{i}\times(-\eta^{-1},\eta^{-1})\rightarrow N_{i}$, $i=1,2$, such that $\psi_{1}(\mathbb{S}^{n-1}/\Gamma_{1}\times(-\eta^{-1}+2,\eta^{-1}-2))\cap\psi_{2}(\mathbb{S}^{n-1}/\Gamma_{2}\times(-\eta^{-1}+1,\eta^{-1}-1))\neq\emptyset$. Then $\Gamma_{1}$ and $\Gamma_{2}$ are conjugate in $O(n)$. Moreover, if $\psi_{1}(\mathbb{S}^{n-1}/\Gamma_{1}\times(-\eta^{-1},-\eta^{-1}+0.01))\cap N_{2}=\emptyset$ and $\psi_{2}(\mathbb{S}^{n-1}/\Gamma_{2}\times(\eta^{-1}-0.01,\eta^{-1}))\cap N_{1}=\emptyset$, there exists a Hamilton’s canonical parametrization $\phi:\mathbb{S}^{n-1}/\Gamma_{1}\times(a,b)\rightarrow\mathcal{T}\subset N_{1}\cup N_{2}$ which gives an $\varepsilon$-tube $\mathcal{T}$ in $N_{1}\cup N_{2}$ with $\psi_{1}(\mathbb{S}^{n-1}/\Gamma_{1}\times(-\eta^{-1}+1,\eta^{-1}))\cup\psi_{2}(\mathbb{S}^{n-1}/\Gamma_{2}\times(-\eta^{-1},\eta^{-1}-1))\subset\mathcal{T}$.
@@
Proof.   Given $0<\varepsilon\leq\varepsilon_{2}$ and $\eta$-neck $N_{i}$ ($i=1,2$) in $(\mathcal{O},g)$ with $\eta\leq\tilde{\varepsilon}_{1}(\varepsilon)$, by Lemma A.6, there exists a Hamilton’s canonical parametrization $\phi_{i}:\mathbb{S}^{n-1}/\Gamma_{i}\times(a_{i},b_{i})\rightarrow\mathcal{T}_{i}\subset N_{i}$ which gives an $\varepsilon$-tube $\mathcal{T}_{i}$ in $N_{i}$ with $\psi_{i}(\mathbb{S}^{n-1}/\Gamma_{i}\times(-\eta^{-1}+1,\eta^{-1}-1))\subset\mathcal{T}_{i}$. We can take $\phi_{i}$ to be maximal in the sense that $\mathcal{T}_{i}$ cannot be properly included in the image of any Hamilton’s canonical parametrization which gives an $\varepsilon$-tube in $N_{i}$, $i=1,2$. Let $\Phi_{i}=\phi_{i}\circ\pi_{i}:\mathbb{S}^{n-1}\times(a_{i},b_{i})\rightarrow\mathcal{T}_{i}\subset N_{i}$ be Hamilton’s canonical uniformization of $\mathcal{T}_{i}$, where $\pi_{i}:\mathbb{S}^{n-1}\times(a_{i},b_{i})\rightarrow\mathbb{S}^{n-1}/\Gamma_{i}\times(a_{i},b_{i})$ is the natural projection. Choose a point $x_{0}\in\psi_{1}(\mathbb{S}^{n-1}/\Gamma_{1}\times(-\eta^{-1}+2,\eta^{-1}-2))\cap\psi_{2}(\mathbb{S}^{n-1}/\Gamma_{2}\times(-\eta^{-1}+1,\eta^{-1}-1))$. Then there exists $s_{1}\in(a_{1},b_{1})$ away from $a_{1}$ and $b_{1}$, and $s_{2}\in(a_{2},b_{2})$, such that $x_{0}\in\Phi_{1}(\mathbb{S}^{n-1}\times\{s_{1}\})\cap\Phi_{2}(\mathbb{S}^{n-1}\times\{s_{2}\})$. If $\varepsilon_{3}=\varepsilon_{3}(n)\leq\varepsilon_{2}$ is sufficiently small and $0<\varepsilon\leq\varepsilon_{3}$, by the proof of Theorem C2.4 in [51], which uses Lemma C2.1 in [51], there exists an isometry of the round cylinder $\mathbb{S}^{n-1}\times\mathbb{R}$, say $G$, such that $G(\mathbb{S}^{n-1}\times\{s_{2}\})=\mathbb{S}^{n-1}\times\{s_{1}\}$ and $\Phi_{1}G|_{\mathbb{S}^{n-1}\times\{s_{2}\}}=\Phi_{2}|_{\mathbb{S}^{n-1}\times\{s_{2}\}}$. So $\phi_{1}(\mathbb{S}^{n-1}/\Gamma_{1}\times\{s_{1}\})=\phi_{2}(\mathbb{S}^{n-1}/\Gamma_{2}\times\{s_{2}\})$. It follows that $\Gamma_{1}$ and $\Gamma_{2}$ are conjugate in $O(n)$.
@@
As in Section 72 of [58], there exist numbers $b_{0}$ and $b^{\prime}$ with $\frac{1}{10}<b^{\prime}<b_{0}$, and a smooth convex function $\varphi$ defined on $(-\infty,b_{0}]$ with $\varphi|_{(-\infty,0]}=0$ and $\varphi(z)=ce^{-1/z}$ for $z\in(0,\frac{1}{10}]$, where $c<\frac{c_{n}}{2}$ is a small positive constant, such that $\varphi$ coincides with $\varphi_{k}$ on $[b^{\prime},b_{0}]$ for some $k>0$, and the metric $e^{-2\varphi}g_{\text{cly}}|_{\mathbb{S}^{n-1}\times(-\infty,b_{0}]}$ has nonnegative curvature operator. Here, we use the fact that if $c<\frac{c_{n}}{2}$,